# xTrace: A Facial Expressive Behaviour Analysis Tool for Continuous Affect Recognition

### v1

@@ 3. Video Datasets for Facial Affect Recognition
Figure 3. Comparison of emotion label distributions between the Aff-Wild2 dataset (Kollias et al., 2025)(top) and the xTrace dataset (bottom).
Figure 4. Sample face images from the xTrace VA dataset
Figure 5. VA label distribution in the xTrace training dataset
@@
Limitations of Existing Datasets. SEWA (Kossaifi et al., 2019) and Aff-wild-2 (Kollias and Zafeiriou, 2018) are two prominent in-the-wild video datasets that are used to train and evaluate most existing facial affect recognition models. These datasets enabled numerous advancements in affect recognition (Kollias et al., 2025), especially the Aff-wild-2 dataset through a series of ABAW challenges (Kollias et al., 2020; Kollias and Zafeiriou, 2021). However, they are still relatively small-scale datasets: SEWA contains 538 videos collected from 398 subjects, and Aff-wild-2 is composed of 594 videos from 584 subjects. Considering the complexity of naturalistic affect recognition as a learning task, such small-scale datasets may not be sufficient to cover the variations encountered in naturalistic and in-the-wild emotional expressions. Another key limitation of these datasets is their limited coverage of the Valence and Arousal space. For example, as shown in Figure 3 (a), the label distributions of Aff-wild-2 clearly show insufficient coverage of the affect space. To make the proposed facial affect recognition model in this work robust to real-world variations and cover the 2D emotion space as extensively as possible, we used a well-curated large-scale video dataset constructed from various in-house collected and curated face videos.
@@
Expressive Video Clips Curation. A set of 450k short video clips is curated from various in-house collected and in-the-wild face video datasets. These datasets were sourced from  8.5k subjects. For example, Figure 4 shows some RGB face images sampled from in-the-wild settings. The duration of the extracted short video clips varies in the range of 3 to 5 seconds. This video set also contains a significant proportion of NIR face videos captured in in-lab and driving (in-car) settings. To make sure that the low-arousal quadrants in the emotion space have a representative number of examples in this dataset, video clips sourced from an internal collected drowsiness dataset are also included. This dataset is partitioned into train and validation sets in a subject-independent fashion, resulting in 400k clips for training and 50k clips for validation and testing purposes.
@@
VA Annotation. Considering the prohibitively expensive nature of trace-based per-frame VA annotation approaches, we employed a semi-automated annotation methodology to obtain per-clip VA emotion labels in a scalable approach. In this method, we used a simple web interface tool for single-point VA annotation in which each video clip is labelled with a point in the 2D emotion space. Compared to the label distributions of the Aff-wild-2 datasets, as shown in Figure 3 (b), the overall label distribution of this dataset has much better coverage in the VA space. Representative frames sampled from some example training video clips with different VA values are illustrated in Figure 5.
@@
Inter-Rater Disagreement Analysis. To evaluate the quality of VA labels obtained using the single-point annotation method described above, we performed an inter-rater disagreement analysis. For this purpose, we extracted a subset of 30k video clips annotated by at least three human raters. Since the labels annotated here are not traces but single points, instead of the standard correlation measures, we adopted a simple weighted mean absolute error (WMAE) method for measuring the inter-rater disagreement scores. Here, the weight values used in the MAE function are derived from the inverted distance values between the corresponding points. Given three VA annotations for a video clip, we computed WMAE scores using the weights derived inter-rater reliability scores (Shrout and Fleiss, 1979) calculated for each annotator. On this 30k video subset, the average WMAE values obtained for valence and arousal dimensions are 0.17 and 0.19 respectively. These low WMAE values indicate good inter-rater agreement levels of the annoated VA labels.
4. Evaluation Results
4.1. Affect Feature Descriptor Evaluation
@@
Uncertainty-Aware Landmark Detection Evaluation results are presented in Table 1. For this evaluation, we used an internally curated validation dataset composed of  39k in-the-wild RGB and NIR images. Both metrics, Area Under Curve (AUC) and Normalised Mean Error (NME), clearly show that the landmark detection model in xTrace achieves noticeably better point localisation accuracy despite its low computational complexity. Figure 8 illustrates the ability of landmark-wise uncertainty estimates in detecting occluded facial regions, by comparing visible and occluded landmarks.
RGB Set NIR Set
@@
OpenFace 2.2.0 (Tadas et al., 2018) 7.03 55.46 4.56 60.29
MediaPipe (Lugaresi et al., 2019) 4.36 55.77 6.62 46.69
xTrace 1.59 88.72 1.12 89.68
Table 1. Landmark detection evaluation results
@@
Figure 6. Cumulative Error Distribution (CED) plots of landmark detection models on (a) RGB and (b) NIR images.
@@
AU Intensity Estimation Evaluation results are shown in Table 2, in terms of AU-wise and mean Intra-Class Correlation (ICC) scores. This in-house evaluation set contains  175k RGB and NIR face images annotated with AU intensity values of 15 AUs. Except for two AUs (07 and 14), for all other AUs the ICC scores are above 0.5. Overall, these high ICC scores clearly indicate the strong predictive accuracy of the AU intensity model used in xTrace, and it outperformed OpenFace by 27% in terms of the mean ICC score.
AU 01 02 04 05 06 07 09 10 12 14 15 17 23 25 45 Avg.
OpenFace 2.2.0 0.35 0.27 0.38 0.42 0.74 0.30 0.25 0.22 0.71 0.16 0.57 0.56 0.13 0.59 0.46 0.41
xTrace 0.76 0.70 0.69 0.66 0.75 0.47 0.68 0.53 0.79 0.42 0.76 0.54 0.93 0.78 0.79 0.68
@@
Table 2. Evaluation results of the AU intensity estimation model in xTrace and comparison with OpenFace 2.2.0
Figure 7. AU model accuracy comparison with OpenFace
Figure 8. Landmark uncertainty estimates quality in xTrace
4.2. Affect Recognition Evaluation
Figure 9. xTrace VA evaluation results
@@ 4.2. Affect Recognition Evaluation
Overall VA Accuracy Evaluation Results across the 2D affect space are presented in Figure 9 in terms of two standard evaluation metrics (Kollias et al., 2025; Toisoul et al., 2021): Concordance Correlation Coefficient (CCC) and Mean Absolute Error (MAE). While CCC captures the overall agreement between the model predictions and ground truth labels of all the test samples, MAE is computed between individual (per-frame) predictions and then averaged for all the samples. In total, this internally curated validation set contains  5 million frames from  50k in-the-wild test videos. This validation set is carefully constructed to cover as many bins as possible in the 2D affect space (see the sample distribution maps in Figure 10.b.). The CCC scores and MAE values for both valence and arousal clearly illustrate the strong predictive accuracy of the affect recognition model in xTrace.
(a) MAE distribution map on the VA validation set
(b) Sample distribution map of the VA validation set
Figure 10. Fine-grained VA evaluation results of xTrace
@@
Table 3. VA evaluation with varying head pose angles
Figure 11. xTrace vs. Augsburg Affect Toolbox: (a). overall CCC and (b) and (c). robustness to non-frontal head pose angles.
More fine-grained VA evaluation results (MAE values) are presented in Figure 10, along with the sample distribution map of the evaluation set. As computing CCC scores makes less sense in the case of fine-grained bins, we used MAE scores to compare the model’s accuracy across different VA bins. The bins with MAE values below and above human inter-rater disagreement (Weighted MAE) are marked in green and purple colors respectively. This result clearly illustrates that xTrace performs well in most VA bins, except for a few bins in the low-arousal quadrants. It could be due to the limitations of the visual modality in inferring arousal (Pantic and Rothkrantz, 2000).
VA Accuracy Robustness to Non-Frontal Head Pose is analysed in Table 3 in terms of the CCC scores and the MAE values. The nonfrontal head pose angles considered in this evaluation cover the yaw bins [+60, +30, 0, +30, -60] and the pitch bins [+25, 0, -25]. For a fair comparison, it is ensured that this internal evaluation set is curated for each pose bin so that it has a roughly equal number of video clips ( 3k clips per pose bin). Both CCC and MAE values show that the VA model’s accuracy is highly robust to changes in the head pose angles. It is worth noting that even for the yaw +/- 60 degree range, the VA predictive accuracy drop is very low, compared to the frontal bin’s VA accuracy.
Leave N-in Validation Results of VA Uncertainty estimates are presented in Table 4. In this evaluation protocol, only N% of the validation examples are considered for evaluation based on how high or low their cumulative uncertainty scores are. The CCC and MAE values that correspond to the ‘Lowest Uncertainty Filter’ show how well the model performs when we consider only N% of validation examples with the lowest uncertainty. Similarly, the results of the ‘Highest Uncertainty Filter’ show how worse the model performs when only the examples with the highest uncertainty are considered for evaluation. The trends in the CCC and MAE values in Table 4 for different N values show that the cumulative uncertainty scores are strongly correlated with the predictive accuracy. This strong correlation suggests that we can assess the reliability of xTrace’s affect predictions in downstream applications.
Dimension Metric Filter Type Accuracy Trend (N: 25% $\to$ 100%)
Valence CCC Lowest Uncert. 0.93 $\downarrow$ 0.81
Highest Uncert. 0.61 $\uparrow$ 0.81
MAE Lowest Uncert. 0.08 $\uparrow$ 0.15
Highest Uncert. 0.21 $\downarrow$ 0.15
Arousal CCC Lowest Uncert. 0.94 $\downarrow$ 0.88
Highest Uncert. 0.73 $\uparrow$ 0.88
MAE Lowest Uncert. 0.07 $\uparrow$ 0.14
Highest Uncert. 0.20 $\downarrow$ 0.14
@@
Table 4. Leave-N-in validation results of VA uncertainty
5. Conclusion
In this paper, we introduced xTrace, a robust, explainable and computationally lightweight tool for in-the-wild naturalistic facial expressive behaviour analysis. xTrace is trained on a corpus of 450k face videos, the largest video dataset for affect recognition to date. This makes xTrace capable of handling challenges encountered in real-world settings, ranging from subtle and nuanced expressive cues to partial occlusions caused by non-frontal head pose angles, etc. Some key system design choices we made in building xTrace include using Low-Level Descriptors of facial expressions (for model interpretability and compute efficiency), leveraging long-range temporal context (for temporally consistent affect predictions) and outputting sampling-free uncertainty estimates (for reliability in downstream applications). Through extensive benchmarking of xTrace’s components against existing tools such as MediaPipe, OpenFace and Augsburg Affect Toolbox, we demonstrated its superior accuracy in face alignment, AU intensity estimation and continuous affect recognition tasks. We believe xTrace will accelerate the adoption of facial affect recognition in a wide range of real-world applications supporting emphathetic human-computer interactions.
Safe and Responsible Innovation. The proposed tool for facial expressive behavior analysis has several real-world applications to support empathetic Human-Computer interactions. Towards addressing risks with existing end-to-end AI tools for facial expression analysis, the tool proposed in this work is made more explainable by using High-Level Descriptors of facial expressions. We explicitly recognise that facial expressions are influenced by cultural backgrounds and are not perfect indicators of internal states. Responsible development requires acknowledging these limitations and ensuring data privacy. We commit to promoting ethical guidelines and responsible practices for the application of this tool.

### v2

@@ III Video Datasets for Facial Affect Recognition
Figure 3: Comparison of emotion label distributions between the Aff-Wild2 dataset [11](top) and the xTrace dataset (bottom).
Figure 4: VA label distribution in the xTrace training dataset
@@
Limitations of Existing Datasets. SEWA [15] and Aff-wild-2 [12] are two prominent in-the-wild video datasets that are used to train and evaluate most existing facial affect recognition models. These datasets enabled numerous advances in affect recognition [11], especially the Aff-wild-2 dataset through a series of ABAW challenges [10, 13]. However, they are still relatively small-scale datasets: SEWA contains 538 videos collected from 398 subjects, and Aff-wild-2 is composed of 594 videos from 584 subjects. Considering the complexity of naturalistic affect recognition as a learning task, such small-scale datasets may not be sufficient to cover the variations encountered in naturalistic and in-the-wild emotional expressions. Another key limitation of these datasets is their limited coverage of the valence and arousal space. As shown in Figure 3, the label distributions of Aff-wild-2 clearly show insufficient coverage of the affect space. To make xTrace robust to real-world variations and cover the 2D emotion space as extensively as possible, we used a well-curated large-scale video dataset constructed from various in-house collected and curated face videos.
@@
Expressive Video Clips Curation. A set of $\sim$450k short video clips is curated from various in-house collected and in-the-wild face video datasets. These datasets were sourced from $\sim$8.5k subjects. The duration of the extracted short video clips varies in the range of 3 to 5 seconds. This video set also contains a significant proportion of NIR face videos captured in in-lab and driving (in-car) settings. To make sure that the low-arousal quadrants in the emotion space have a representative number of examples in this dataset, video clips sourced from an internal collected drowsiness dataset are also included. This dataset is partitioned into model development and benchmarking sets in a subject-independent fashion, resulting in $\sim$400k clips for model building and $\sim$50k clips for evaluation purposes. As a large proportion of this data set is composed of in-house data collected from external participants, For a variety of reasons, most notably GDPR, we are not able to make this dataset publicly available. We made sure that all the video data are sourced ethically from in-house data collection activities and other public sources with the required permissions.
It is worth notingh that Aff-wild-2 is available to only the ABAW challenge participants. Also, as Figure 4 shows, the label distribution of Aff-wild-2 is highly biased towards neutral emotion. Hence, we used the in-house constructed benchmarking dataset with $\sim$50k videos for thoroughly evaluating xTrace. Also, we used the SEWA test set for benchmarking xTrace against existing baselines in the literature.
@@
VA Annotation. Considering the prohibitively expensive nature of trace-based per-frame VA annotation approaches, we employed a semi-automated annotation methodology to obtain per-clip VA emotion labels in a scalable approach. In this method, we used a simple web interface tool for single-point VA annotation in which each video clip is labelled with a point in the 2D emotion space. Compared to the label distributions of the Aff-wild-2 datasets, as shown in Figure 3, the overall label distribution of this dataset has much better coverage in the VA space. Representative frames sampled from some example training video clips with different VA values are illustrated in Figure 4.
@@
Inter-Rater Disagreement Analysis. To evaluate the quality of VA labels obtained using the single-point annotation method described above, we performed an inter-rater disagreement analysis. For this purpose, we extracted a subset of $\sim$30k video clips annotated by at least three human raters. Since the labels annotated here are not traces but single points, instead of the standard correlation measures, we adopted a simple weighted mean absolute error (WMAE) method for measuring the inter-rater disagreement scores. Here, the weight values used in the MAE function are derived from the inverted distance values between the corresponding points. Given three VA annotations for a video clip, we computed WMAE scores using the weights derived inter-rater reliability scores [24] calculated for each annotator. On this $\sim$30k video subset, the average WMAE values obtained for valence and arousal dimensions are 0.17 and 0.19 respectively. These low WMAE values indicate good inter-rater agreement levels of the annoated VA labels.
IV Evaluation Results
IV-A Evaluation of Facial Affect Descriptors
@@
Uncertainty-Aware Landmark Detection Evaluation results are presented in Table IV-A. For this evaluation, we used an internally curated benchmarking dataset composed of $\sim$39k in-the-wild RGB and NIR images. Both metrics, Area Under Curve (AUC) and Normalised Mean Error (NME), clearly show that the landmark detection model in xTrace achieves noticeably better point localisation accuracy despite its low computational complexity. Figure 7 illustrates the ability of landmark-wise uncertainty estimates in detecting occluded facial regions, by comparing visible and occluded landmarks.
RGB Test Set NIR Test Set
@@
[…]
MediaPipe [16] 4.36 55.77 $6.62$ $46.69$
[…]
TABLE I: Comparison of facial landmark detection performance. Metrics: Normalised Mean Error (NME) $\downarrow$ - lower is better, and Area Under the Curve (AUC) $\uparrow$ - higher is better.
@@
Figure 5: Cumulative Error Distribution (CED) plots of landmark detection models on (a) RGB and (b) NIR images.
@@
AU Intensity Estimation Evaluation results are shown in Table IV-A, in terms of AU-wise and mean Intra-Class Correlation (ICC) scores. This in-house evaluation set contains $\sim$175k RGB and NIR face images annotated with AU intensity values of 15 AUs. Except for two AUs (07 and 14), for all other AUs the ICC scores are above 0.5. Overall, these high ICC scores show the strong predictive accuracy of the AU intensity model used in xTrace, and it outperformed OpenFace by $\sim$27% in terms of the mean ICC score.
AU $\rightarrow$ 01 02 04 05 06 07 09 10 12 14 15 17 23 25 45 Avg.
[…]
@@
TABLE II: Evaluation results of the AU intensity estimation model in xTrace and comparison with OpenFace 2.2.0
Figure 6: AU model accuracy comparison with OpenFace
Figure 7: Landmark uncertainty estimates quality in xTrace
IV-B Affect Recognition Evaluation
Method Vale. CCC Arou. CCC Avg. CCC
Augs. Aff. Toolbox [17] $0.62$ $0.11$ $0.37$
Mitekova et al. [18] $0.44$ $0.39$ $0.42$
Toisoul et al. [30] $0.65$ $0.61$ $0.63$
Kossafi et al. [14] $0.75$ $0.52$ $0.64$
ResNet-Aff. Proc. [27] $0.73$ $0.58$ $0.66$
HGFAN-Aff. Proc. [27] $0.74$ $0.63$ $0.69$
Affective Processes [23] $0.75$ $0.64$ $0.70$
[…]
TABLE III: VA benchmarking results on the SEWA test set.
Figure 8: Emotion quadrant-wise VA evaluation results
@@ IV-B Affect Recognition Evaluation
Overall VA Accuracy Evaluation Results across the 2D affect space are presented in Figure 8 in terms of two standard evaluation metrics [11, 30]: Concordance Correlation Coefficient (CCC) and Mean Absolute Error (MAE). While CCC captures the overall agreement between the model predictions and ground truth labels of all the test samples, MAE is computed between individual (per-frame) predictions and then averaged for all the samples. In total, this internally curated benchmarking set contains $\sim$5 million frames from $\sim$50k in-the-wild test videos. This benchmarking set is carefully constructed to cover as many bins as possible in the 2D affect space (see the sample distribution maps in Figure 9.b.). The CCC scores and MAE values for both valence and arousal illustrate the strong predictive accuracy of the affect recognition model in xTrace.
SEWA Benchmarking Results are presented in Table III. This test set contains 49342 frames labelled with valence and arousal, sourced from 53 in-the-wild videos of subjects from different cultures. This evaluation clearly shows that xTrace outperforms strong existing baselines such as Affective Processes [23] by a large margin (7.1% relative improvement in mean CCC). In addition, xTrace performs substantially better than the Augsburg Affect Toolbox, almost twofold. These results illustrate the superior performance of xTrace on a standard facial affect recognition benchmarking dataset.
(a) MAE distribution on the VA benchmarking set
(b) Sample distribution of the VA benchmarking set
Figure 9: Fine-grained VA evaluation results
Figure 10: xTrace vs. Augsburg Affect Toolbox: (a). overall CCC and (b) and (c). robustness to non-frontal head pose angles.
More fine-grained VA evaluation results (MAE values) are presented in Figure 9, along with the sample distribution map of the evaluation set. Here, we used MAE scores to compare the model’s accuracy across different VA bins. The bins with MAE values below and above human inter-rater disagreement (Weighted MAE) are marked in green and purple colors respectively. This result clearly illustrates that xTrace performs well in most VA bins, except for a few bins in the low-arousal quadrants. It could be due to the limitations of the visual modality in inferring arousal [20].
@@
TABLE IV: VA evaluation with varying head pose angles
VA Accuracy Robustness to Non-Frontal Head Pose is analysed in Table IV in terms of the CCC scores and the MAE values. The nonfrontal head pose angles considered in this evaluation cover the yaw bins [+60, +30, 0, +30, -60] and the pitch bins [+25, 0, -25]. For a fair comparison, it is ensured that this internal evaluation set is curated for each pose bin so that it has a roughly equal number of video clips ($\sim$3k clips per pose bin). Both CCC and MAE values show that the VA model’s accuracy is highly robust to changes in the head pose angles. It is worth noting that even for the yaw +/- 60 degree range, the VA predictive accuracy drop is very low, compared to the frontal bin’s VA accuracy.
Leave N-in Validation Results of VA Uncertainty estimates are presented in Table V. In this evaluation protocol, only N% of the test examples are considered for evaluation based on how high or low their cumulative uncertainty scores are. The CCC and MAE values that correspond to the ‘Lowest Uncertainty Filter’ show how well the model performs when we consider only N% of test examples with the lowest uncertainty. Similarly, the results of the ‘Highest Uncertainty Filter’ show how worse the model performs when only the examples with the highest uncertainty are considered for evaluation. The trends in the CCC and MAE values in Table V for different N values show that the cumulative uncertainty scores are strongly correlated with accuracy. This strong correlation suggests that we can assess the reliability of xTrace’s affect predictions in downstream applications.
TABLE V: Leave-N-in validation results of VA uncertainty
@@
Valence CCC Lowest Unct. 0.93 $\downarrow$ 0.81
Lowest Unct. 0.61 $\uparrow$ 0.81
MAE Lowest Unct. 0.08 $\uparrow$ 0.15
Lowest Unct. 0.21 $\downarrow$ 0.15
Arousal CCC Lowest Unct. 0.94 $\downarrow$ 0.88
Lowest Unct. 0.73 $\uparrow$ 0.88
MAE Lowest Unct. 0.07 $\uparrow$ 0.14
Lowest Unct. 0.20 $\downarrow$ 0.14
Accuracy Trend: Green $\uparrow\downarrow$ - Improving accuracy, Red $\uparrow\downarrow$ - Degrading accuracy
V Conclusion
In this paper, we introduced xTrace, a robust, explainable and computationally lightweight tool for in-the-wild naturalistic facial expressive behaviour analysis. xTrace is trained on a corpus of $\sim$450k face videos, the largest video dataset for affect recognition to date. This makes xTrace capable of handling challenges encountered in real-world settings, ranging from subtle and nuanced expressive cues to partial occlusions caused by non-frontal head pose angles, etc. Some key system design choices we made in building xTrace include using Low-Level Descriptors of facial expressions (for model interpretability and compute efficiency) and outputting sampling-free uncertainty estimates (for reliability in downstream applications). Through extensive benchmarking of xTrace’s components against existing tools such as MediaPipe, OpenFace and Augsburg Affect Toolbox, we demonstrated its superior accuracy in face alignment, AU intensity estimation and continuous affect recognition tasks. We believe xTrace will accelerate the adoption of facial affect recognition in a wide range of real-world applications supporting emphathetic human-computer interactions.
Ethical Impact Statement. xTrace has several real-world applications to support empathetic Human-Computer interactions. We commit to promoting ethical guidelines and responsible practices for the application of this tool. The intended use cases of xTrace include, but are not limited to, mobile applications for mental health and wellbeing, social robotics, driver monitoring systems, etc. It is heavily optimised for on-device video processing, which is essential to ensure data privacy in the aforementioned use cases.